OpenStudy (study_buddy99):

How can I factor 27x^3+8?

OpenStudy (alexandervonhumboldt2):

hint: \[27x^3+8=3^3x^3+2^3\] use the sum of cubes formula: \[a^3+b^3=(a+b)*(a^2-ab+b^2)\]

OpenStudy (study_buddy99):

I'm still confused...

OpenStudy (landon34):

:/

OpenStudy (study_buddy99):

/: